OpenStudy (anonymous):

According to the natural sequence of exponents, how would you explain 2^-3

OpenStudy (anonymous):

Please help, I am confused.

OpenStudy (anonymous):

2^4 is 16. Each time the exponent decreases by one, the value of the expression is divided by 2. So 2^3=8, 2^2=4, 2^1=2, 2^0=1, 2^-1=1/2, 2^-2=1/2, and 2^-3=1/8

OpenStudy (anonymous):

I do not understand how it is decreased by 1, 2^0 is 1 and 2^-1 is 1/2 that is decreased by .5, not one.

OpenStudy (anonymous):

Let we have a base 2 with an exponent power raised to -3 So in order to have a positive exponent factor, we take reciprocal of the number, 1/2^3 Now that the base has a positive exponent number, we simply multiply it the times on the exponent number, so: 1/2*2*2=1/8 is our answer.

OpenStudy (anonymous):

I think I understand, thank you.

OpenStudy (anonymous):

your welcome